# A Lightweight Video Anomaly Detection Model with Weak Supervision and Adaptive Instance Selection

### v1

@@ III-F Antagonistic Loss Function
$\begin{split}loss_{antagonistic}=\left[1-(S_{top-1}^{P}-S_{top-1}^{N})\right]\\ +S_{top-1}^{N}+(1-S_{top-1}^{P})\end{split}$ (8)
Here, $S_{top-1}^{P}$ and $S_{top-1}^{N}$ respectively represent the highest (top-1) abnormality score of instances in the positive and negative bags predicted by the model. $loss_{antagonistic}$ not only embodies the antagonistic constraint between positive and negative instances, but also provides optimization constraints for them respectively. Finally, the total loss function of our method Light-WSVAD is as follows:
@@
IV Performance Evaluationr
IV-A Datasets and Evaluation Metrics
@@ IV-B Compared Methods
We compare our method with a number of existing works, including twelve general (non-lightweight) models and two lightweight methods. For some models, we consider two versions of using different feature extractors.
@@
Among the twelve non-lightweight models, the work of Sultani et al. [3] is the first to use the MIL framework and weakly supervision for video anomaly detection, and most of the subsequent works follow their settings. The method of Zhang et al. [4] defines a new inner-bag loss in the MIL framework to limit the function space and expand the differences in feature distributions for different types of instances. ARNet [6] uses a dynamic multi-instance learning loss to enlarge the interclass distance between anomalous and normal instances, and a center loss to narrow the intraclass distance of normal instances to boost the model’s ability to distinguish anomalies. CLAWS Net+ [41] employs a clustering loss to mitigate labeling noise, which improves the learning of representations for both abnormal and normal videos. MIST [36] adopts a multi-instance self-training framework to effectively refine task-specific discriminative representations using only video-level annotations. Similarly, MSL [20] proposes a self-training strategy and designs a transformer-based Multi-Sequence Learning network to further learn video-level anomaly probabilities and clip-level anomaly scores. RTFM [7] trains the feature magnitude learning function to effectively identify positive instances and improves the robustness of the model. BN-SVP [42] introduces a novel approach called Bayesian non-parametric submodular video partitioning to enhance the training of MIL models, and provides a reliable solution for robust anomaly detection in practical scenarios that involve outlier instances or multiple types of abnormal events. DAR [39] and the method of Wu et al. [5] design the model structure using a multi-branching and multi-stage approach, respectively, to improve the model’s ability to understand and fuse different modal data in the videos. Mu et al.’s method  [43] and GCN-Anomaly  [22] innovatively employ graph convolutional networks for enhancing the model’s ability to understand spatio-temporal correlation features in videos. Despite their remarkable performance in anomaly detection, these models do not consider factors such as model parameter size and runtime speed in their designs.
In the two lightweight methods, the work of Chang et al.[8] proposes a lightweight MIL model incorporating a comparative attention module to improve model performance. The method of [9] introduces a self-attention mechanism to enhance the performance of the lightweight MIL model. Although these two methods achieve considerable advance, there remains substantial room to enhance the design of the model and boost the performance.
IV-C Implementation Details
@@
IV-D Performance Comparison
To demonstrate our method’s effectiveness, we compare it with existing anomaly detection methods, including general WSVAD models and lightweight WSVAD methods.
Table I presents the frame-level AUC results on the ShanghaiTech dataset. We can see that our method obtains the highest performance of 95.9% among the lightweight weakly-supervised methods, and surpasses most of general (non-lightweight) methods, whilst maintaining a parameter size of merely 0.14M.
Table II gives the frame-level AUC performance on the UCF dataset. Our method not only achieves the highest AUC of 84.7% among the lightweight weakly-supervised methods, but also outperforms most non-lightweight weakly-supervised methods, standing at the third place among all general models.
TABLE I: Performance comparison of frame-level AUC on ShanghaiTech. For lightweight models, the best result is in red and the second best in blue; For general (Non-lightweight) models, the best result is in bold.
@@ IV-D Performance Comparison
Lightweight Models Light-WSVAD (ours) I3D 95.9
Besides comparing the AUC results of different methods on the two datasets, we also compare the number of parameters among all methods, the results are presented in Table III. We can see that our method utilizes only 0.14M parameters, which is the lowest among the compared models. Particularly, our method has only less than 1% of the parameters of the RTMF method, and only half of the parameters of the currently smallest resource-intensive lightweight method [8].
Combined the comparison results from Table I to Table III, it is obvious that our method is a lightweight yet good-performance method for weakly supervised video anomaly detection.
TABLE II: Performance comparison of frame-level AUC on UCF. For lightweight models, the best result is in red and the second best in blue; For general (Non-lightweight) models, the best result is in bold.
@@
Lightweight Models Light-WSVAD (ours) I3D 84.7
TABLE III: Model size comparison among weakly supervised methods. The best result is in red and the second best is in blue.
@@
Light-WSVAD (ours) 0.14
IV-E Ablation Studies
Effect of major modules in Light-WSVAD. To evaluate the effectiveness of the major modules in our method, we conduct ablation experiment on the ShanghaiTech dataset. The results are presented in Table IV. Here, the baseline model is a simple network consisting of fully connected layers, with an AUC of 93.8%. The results in Table IV show that employing MTA, AIS, and the antagonistic loss (A-Loss in short) to the baseline individually can obviously improve the model’s performance. Furthermore, our proposed lightweight HFC structure does not have negative impact on the model’s performance. In summary, by combining MTA, HFC, AIS, and A-Loss into our method, a 2.1% performance improvement on the ShanghaiTech dataset is achieved.
TABLE IV: Ablation study on ShanghaiTech.
@@ IV-E Ablation Studies
Effect of parameter $k$ in MTA. MTA is to capture multi-level temporal correlations of consecutive instances by integrating inter-instance feature relationships across multiple time intervals. As described in Section III-C, MTA has a hyperparameter $k$, which represents the maximum number of consecutive $k$ instances, from which MTA can extract temporal correlation information. The value of $k$ affects the model’s performance. We change the value of $k$ from 3 to 15 with a stepsize of 2, and report the performance results in Table V. We can see that when $k$ is set to 5, the model achieves the best performance. By analyzing the performance change when setting different values of k for MTA, we can see that abnormal events typically have a relatively short duration. When $k$ is too large, the proportion of normal instances is too high, which may dampen the influence of abnormal instances so that MTA fails to provide discriminative temporal correlation information. On the other hand, a too small $k$ may result in insufficient coverage of anomalous instances, so that MTA cannot obtain effective temporal correlation information by enough inter-instance information.
TABLE V: Ablation study on parameter $k$ in MTA on ShanghaiTech.
@@
Effect of the antagonistic loss function. To further verify the advantage of the proposed antagonistic loss, we compare the performance of three configurations: not using the sparsity loss, using the sparsity loss, and using our antagonistic loss (our method) on the ShanghaiTeah dataset, the results are presented in Table VI. We can see that the antagonistic loss can obviously improve model performance, whereas the sparsity loss results in a decrease in model performance.
@@
TABLE VI: Ablation study on loss functions in our method on ShanghaiTech.
@@ IV-F Visual Analysis
Figure 6: Visualization of test results of our method Light-WSVAD and the baseline model on the ShanghaiTech dataset. The pink area denotes the time period that anomalies occur in the video, and the numerical labels on the line correspond to the labels in the video frames. In the anomalous video screenshots, the numbers are highlighted in red, and the anomalous objects are marked with red boxes in the frames.
Figure 7: Visualization of test results of our method Light-WSVAD and the baseline model on the UCF-Crime dataset. The pink area denotes the time period that anomalies occur in the video, and the numerical labels on the line correspond to the labels in the video frames. In the anomalous video screenshots, the numbers are highlighted in red, and the anomalous objects are marked with red boxes in the frames.
@@
V Conclusion
This study develops a lightweight weakly-supervised video anomaly detection method (Light-WSVAD) that can effectively addresses the uncertainty and high-parameter issues associated with the existing WSVAD methods. Compared with existing lightweight models, Light-WSVAD has the smallest number of parameters and the best performance. However, predicting sudden anomalies without warning signals is still a serious challenge for current video anomaly detection methods. And the predicted results on different anomalous classes in the UCF-Crime dataset also show that our model is not effective enough in detecting sudden anomaly classes. In the future, we plan to employ multi-modal techniques to enhance the model’s understanding of video content, and thus develop more effective video anomaly detection models.

### v2

@@ 3.6 Antagonistic Loss Function
$\begin{split}loss_{antagonistic}=S_{top-1}^{N}+(1-S_{top-1}^{P})\end{split}$ (8)
Here, $S_{top-1}^{P}$ and $S_{top-1}^{N}$ respectively represent the highest (top-1) abnormality score of instances in the positive and negative bags predicted by the model. $loss_{antagonistic}$ not only embodies the antagonistic constraint between positive and negative instances, but also provides optimization constraints for them respectively. Finally, the total loss function of our method Light-WVAD is as follows:
@@
4 Performance Evaluation
4.1 Datasets and Evaluation Metrics
@@ 4.2 Compared Methods
We compare our method with a number of existing works, including thirteen general (non-lightweight) models and there lightweight methods. For some models, we consider two versions of using different feature extractors.
@@
Among the thirteen non-lightweight models, the work of Sultani et al. [3] is the first to use the MIL framework and weakly supervision for video anomaly detection, and most of the subsequent works follow their settings. The method of Zhang et al. [10] defines a new inner-bag loss in the MIL framework to limit the function space and expand the differences in feature distributions for different types of instances. ARNet [12] uses a dynamic multi-instance learning loss to enlarge the interclass distance between anomalous and normal instances, and a center loss to narrow the intraclass distance of normal instances to boost the model’s ability to distinguish anomalies. CLAWS Net+ [48] employs a clustering loss to mitigate labeling noise, which improves the learning of representations for both abnormal and normal videos. MIST [43] adopts a multi-instance self-training framework to effectively refine task-specific discriminative representations using only video-level annotations. Similarly, MSL [27] proposes a self-training strategy and designs a transformer-based Multi-Sequence Learning network to further learn video-level anomaly probabilities and clip-level anomaly scores. RTFM [13] trains the feature magnitude learning function to effectively identify positive instances and improves the robustness of the model. BN-SVP [4] introduces a novel approach called Bayesian non-parametric submodular video partitioning to enhance the training of MIL models, and provides a reliable solution for robust anomaly detection in practical scenarios that involve outlier instances or multiple types of abnormal events. DAR [46] and the method of Wu et al. [11] design the model structure using a multi-branching and multi-stage approach, respectively, to improve the model’s ability to understand and fuse different modal data in the videos. Mu et al.’s method  [49] and GCN-Anomaly  [29] innovatively employ graph convolutional networks for enhancing the model’s ability to understand spatio-temporal correlation features in videos. The NTCN-ML [51] model extracts temporal representations of video data to construct a time-series pattern to optimize the multi-instance learning process. MGFN [52] propose a novel glance and focus network to effectively integrate spatial-temporal information for accurate anomaly detection.In addition, MGFN [52] propose the Feature Amplifcation Mechanism and a Magnitude Contrastive Loss to enhance the discriminativeness of feature magnitudes for detecting anomalies. NG-MIL [53] encodes diverse normal patterns from noise-free normal videos into prototypes to construct a similarity-based classifier. By combining predictions from classifiers, this approach can refine anomaly scores, reducing training instability from weak labels. CU-Net [54] introduces an enhanced framework with a two-stage self-supervised model. This model can generate and iteratively refine pseudo-labels by leveraging the completeness and uncertainty properties of weakly labeled data. HSN [55] propose a Human-Scene Network to learn discriminative representations by capturing both subtle and strong cues in a dissociative manner. Despite their remarkable performance in anomaly detection, these models do not consider factors such as model parameter size and runtime speed in their designs. In the there lightweight methods, the work of Chang et al.[14] proposes a lightweight MIL model incorporating a comparative attention module to improve model performance. The method of [15] introduces a self-attention mechanism to enhance the performance of the lightweight MIL model. BE-WVAD [16] proposes short-input inference modes, which can significantly reduce the required length of input videos and therefore greatly cut down memory and computational costs. Although these two methods achieve considerable advance, there remains substantial room to enhance the design of the model and boost the performance.
4.3 Implementation Details
@@
4.4 Performance Comparison
To demonstrate our method’s effectiveness, we compare it with existing anomaly detection methods, including general WVAD models and lightweight WVAD methods.
Table 1 presents the frame-level AUC results on the ShanghaiTech dataset. We can see that our method obtains the highest performance of 95.9% among the lightweight weakly-supervised methods, and surpasses most of general (non-lightweight) methods, whilst maintaining a parameter size of merely 0.14M.
Table 2 gives the frame-level AUC performance on the UCF dataset. Our method not only achieves the highest AUC of 84.7% among the lightweight weakly-supervised methods, but also outperforms most non-lightweight weakly-supervised methods, standing at the third place among all general models.
Table 3 gives the frame-level AUC performance on the XD-Violence dataset. Our method not only achieves the highest AUC of 77.3% among the lightweight weakly-supervised methods, but also performs comparably to most of the non-lightweight weakly-supervised methods.
Table 1: Performance comparison of frame-level AUC on ShanghaiTech. For lightweight models, the best result is in red and the second best in blue; For general (Non-lightweight) models, the best result is in bold.
@@ 4.4 Performance Comparison
BE-WVAD [16] I3D 95.0
Lightweight Models Light-WVAD (ours) I3D 95.9
Besides comparing the AUC results of different methods on the two datasets, we also compare the number of parameters among all methods, the results are presented in Table 4. We can see that our method utilizes only 0.14M parameters, which is the lowest among the compared models. Particularly, our method has only less than 1% of the parameters of the RTMF method, and only half of the parameters of the currently smallest resource-intensive lightweight method [14].
Combined the comparison results from Table 1 to Table 4, it is obvious that our method is a lightweight yet good-performance method for weakly supervised video anomaly detection.
Table 2: Performance comparison of frame-level AUC on UCF. For lightweight models, the best result is in red and the second best in blue; For general (Non-lightweight) models, the best result is in bold.
@@
BE-WVAD. [16] I3D 84.1
Lightweight Models Light-WVAD (ours) I3D 84.7
Table 3: Performance comparison of frame-level AP on XD-Violence. For lightweight models, the best result is in red and the second best in blue; For general (Non-lightweight) models, the best result is in bold.
Model type Method Feature extractor AP(%)
General Models Sultani et al. [3] C3D 73.2
Wu et al. [11] I3D 75.4
RTFM [13] C3D 75.9
RTFM [13] I3D 77.8
MSL [27] C3D 75.5
MSL [27] I3D 78.3
DAR [46] I3D 78.9
MGFN [52] I3D 79.2
NG-MIL [53] I3D 78.5
CU-Net [54] I3D 78.7
Lightweight Models Chang et al. [14] I3D+flow 71.5
Chang et al. [14] I3D 76.9
BE-WVAD [16] I3D 74.9
Light-WVAD(ours) I3D 77.3
Table 4: Model size comparison among weakly supervised methods. The best result is in red and the second best is in blue.
@@
Light-WVAD (ours) 0.14
4.5 Ablation Studies
Effect of major modules in Light-WVAD. To evaluate the effectiveness of the major modules in our method, we conduct ablation experiment on the ShanghaiTech dataset. The results are presented in Table 5. Here, the baseline model is a simple network consisting of fully connected layers, with an AUC of 93.8%. The results in Table 5 show that employing MTA, AIS, and the antagonistic loss (A-Loss in short) to the baseline individually can obviously improve the model’s performance. Furthermore, our proposed lightweight HFC structure does not have negative impact on the model’s performance. In summary, by combining MTA, HFC, AIS, and A-Loss into our method, a 2.1% performance improvement on the ShanghaiTech dataset is achieved.
Table 5: Ablation study on ShanghaiTech.
@@ 4.5 Ablation Studies
Effect of parameter $k$ in MTA. MTA is to capture multi-level temporal correlations of consecutive instances by integrating inter-instance feature relationships across multiple time intervals. As described in Section 3.3, MTA has a hyperparameter $k$, which represents the maximum number of consecutive $k$ instances, from which MTA can extract temporal correlation information. The value of $k$ affects the model’s performance. We change the value of $k$ from 3 to 15 with a stepsize of 2, and report the performance results in Table 6. We can see that when $k$ is set to 5, the model achieves the best performance. By analyzing the performance change when setting different values of k for MTA, we can see that abnormal events typically have a relatively short duration. When $k$ is too large, the proportion of normal instances is too high, which may dampen the influence of abnormal instances so that MTA fails to provide discriminative temporal correlation information. On the other hand, a too small $k$ may result in insufficient coverage of anomalous instances, so that MTA cannot obtain effective temporal correlation information by enough inter-instance information.
Table 6: Ablation study on parameter $k$ in MTA on ShanghaiTech.
@@
Effect of the antagonistic loss function. To further verify the advantage of the proposed antagonistic loss, we compare the performance of three configurations: not using the sparsity loss, using the sparsity loss, and using our antagonistic loss (our method) on the ShanghaiTeah dataset, the results are presented in Table 7. We can see that the antagonistic loss can obviously improve model performance, whereas the sparsity loss results in a decrease in model performance.
@@
Table 7: Ablation study on loss functions in our method on ShanghaiTech.
@@ 4.6 Visual Analysis
Figure 6: Visualization of test results of our method Light-WVAD and the baseline model on the ShanghaiTech dataset. The pink area denotes the time period that anomalies occur in the video, and the numerical labels on the line correspond to the labels in the video frames. In the anomalous video screenshots, the numbers are highlighted in red, and the anomalous objects are marked with red boxes in the frames.
Figure 7: Visualization of test results of our method Light-WVAD and the baseline model on the UCF-Crime dataset. The pink area denotes the time period that anomalies occur in the video, and the numerical labels on the line correspond to the labels in the video frames. In the anomalous video screenshots, the numbers are highlighted in red, and the anomalous objects are marked with red boxes in the frames.
@@
5 Conclusion
This study develops a lightweight weakly-supervised video anomaly detection method (Light-WVAD) that can effectively addresses the uncertainty and high-parameter issues associated with the existing WVAD methods. Compared with existing lightweight models, Light-WVAD has the smallest number of parameters and the best performance. However, predicting sudden anomalies without warning signals is still a serious challenge for current video anomaly detection methods. And the predicted results on different anomalous classes in the UCF-Crime dataset also show that our model is not effective enough in detecting sudden anomaly classes. In the future, we plan to employ multi-modal techniques to enhance the model’s understanding of video content, and thus develop more effective video anomaly detection models.
[…]
Acknowledgement
This work was supported in part by National Natural Science Foundation of China (No. U1936205, No. 6217230 0), National Key R&D Program of China (No. 2021YFC330030 0), Open Research Projects of Zhejiang Lab (No. 2021KH0A B04), and the Fundamental Research Funds for the Central Universities (No. ZD-21-202101). Finally, we would like to thank Ang Li and Jun Yan for their corrections to the writing of this paper.